As a policy of quick serving restaurant. A customer complaint needs to be attended immediately. On an average on a given day 4 customers complaints. Find the following probabilities.

A) P ( No complaints)

B) P ( At most 2 complaints)


1
Expert's answer
2021-02-24T07:36:17-0500

Poisson (X)=per 4 customer

the required probability "P(x)=\frac{e^{-\mu}\times \mu^n}{x!}"

x>1 so

"P(0)=\frac{e^{-4}\times 4^0}{1}=\frac{1}{e^4}"

Now, n=2

"P(x)=\frac{e^{-4}\times 4^2}{0!}=\frac{16}{e^{4}}"